# Decomposition of Borel graphs and cohomology

### v2

@@ 4.3 Structure of $\textup{H}^{1}(G,R_{G})$
$\sigma(a)(x)=\sum_{y\in[x]_{G}}a(x,y)\sigma\left(1_{\{(x,y)\}}\right)(x).$ (4.5)
@@
$\sigma(a)(x)$ $=\left(1_{\{x\}}\sigma(a)\right)(x)=\sigma\left(1_{\{x\}}a\right)(x)$
@@
$\Psi_{\sigma}(y,x)=\sigma\left(1_{\{(x,y)\}}\right)(x)$
@@
$\Phi_{\Psi_{\sigma}}(a)(x)=\sum_{y\in[x]_{G}}a(x,y)\Psi_{\sigma}(y,x)=\sum_{y\in[x]_{G}}a(x,y)\sigma\left(1_{\{(x,y)\}}\right)(x)=\sigma(a)(x),$
@@
$\cdots\to P_{n}\to P_{n-1}\to\cdots\to P_{0}\to l^{\infty}_{R}(X)\to 0.$ (4.6)
@@
In other words, the set $Z_{G}^{k}$ consists of $\xi\in l^{\infty}_{R}(E_{G})$ such that the function $\xi(\cdot,x)$ is constant on each $G$-connected component of $[x]_{G}\setminus B_{G}(k;x)$. Set $Z_{G}=\bigcup_{k=0}^{\infty}Z_{G}^{k}$. For $\xi\in Z_{G}$ and $a\in R_{G}$, define
@@
This function belongs to $Z_{G}$. Indeed, if $\xi\in Z_{G}^{k}$ and $\varphi\in\mathopen{\hbox{${\langle}$}\kern-1.94444pt\leavevmode\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\leavevmode\hbox{${\rangle}$}}$⟩ with $\textup{graph}\varphi\subset G^{l}$, then we have $(\xi\alpha_{\varphi})(y,x)=\xi(y,\varphi x)1_{\textup{dom}\varphi}(x)$, and thus $\xi\alpha_{\varphi}\in Z_{G}^{k+l}$.
@@
Take $\{\varphi_{i}\}_{i=1}^{n}\subset\mathopen{\hbox{${\langle}$}\kern-1.94444pt\leavevmode\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\leavevmode\hbox{${\rangle}$}}$⟩ so that $G=\bigsqcup_{i=1}^{n}\left(\textup{graph}\varphi_{i}\sqcup\textup{graph}\varphi_{i}^{-1}\right)$. Then by Corollary 4.7, $\textup{H}^{0}(G,W_{G})$ is identified with the right $R_{G}$-submodule
@@
$((\alpha_{\varphi_{i}}-1_{\textup{im}\varphi_{i}})\xi)(y,x)=1_{\textup{im}\varphi_{i}}(y)\left(\xi\left(\varphi_{i}^{-1}y,x\right)-\xi(y,x)\right).$
@@
$y\in[x]_{G}\mapsto 1_{\textup{im}\varphi_{i}}(y)\left(\xi\left(\varphi_{i}^{-1}y,x\right)-\xi(y,x)\right)$
@@
[…]
@@
It suffices to show that for every $k\in\mathbb{Z}_{\geq 0}$, the set $Z_{G}^{k}$ is contained in a finitely generated submodule of $Z_{G}$. Fix $k\in\mathbb{Z}_{\geq 0}$. For $x\in X$, let $\{C_{i}(x)\}_{i=1}^{n}$ be a family of $G$-connected components of $[x]_{G}\setminus B_{G}(k;x)$ such that $[x]_{G}\setminus B_{G}(k;x)=\bigsqcup_{i=1}^{n}C_{i}(x)$. By allowing $C_{i}(x)=\varnothing$, we may assume that the number $n$ does not depend on $x$. Now set $\xi_{i}(\cdot,x)=1_{C_{i}(x)}$ for $x\in X$ and $1\leq i\leq n$. Then for every $\xi\in Z_{G}^{k}$, there exists $\{f_{i}\}_{i=1}^{n}\subset l^{\infty}_{R}(X)$ such that $\xi-\sum_{i=1}^{n}\xi_{i}f_{i}\in R_{G}^{k}$. Hence $Z_{G}^{k}$ is contained in the submodule of $Z_{G}$ genetated by $\{\xi_{i}\}_{i=1}^{n}\cup\{1_{X}\}$. ∎
4.4 Cutsets and $\textup{H}^{1}(G,R_{G})$
@@
(i) Suppose that $\textup{H}^{1}(G,R_{G})$ is finitely generated as a right $R_{G}$-module. Then by Corollary 4.20, $Z_{G}$ is generated by $Z_{G}^{k}$ for some $k$. Let $[x]_{G}\setminus B_{G}(k;x)=\bigsqcup_{i=1}^{n}C_{i}(x)$ as in the proof of Corollary 4.20. Then every element of $Z_{G}^{k}$ is represented by the cutset
@@
which has uniformly bounded boundaries. Note that if $\xi\in Z_{G}$ satisfies $\xi(\cdot,x)=1_{C(x)}$, then for $\varphi\in\mathopen{\hbox{${\langle}$}\kern-1.94444pt\leavevmode\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\leavevmode\hbox{${\rangle}$}}$⟩, we have $(\xi\alpha_{\varphi})(\cdot,x)=1_{C(\varphi x)}1_{\textup{dom}\varphi}(x)$. Hence every element of $Z_{G}$ is represented by the cutset $\mathcal{C}$.
@@
Let $\xi\in Z_{G}$ be such that $\xi(\cdot,x)=1_{C(x)}$ with $C(x)\in\mathcal{C}\cup\{\varnothing\}$. It suffices to show that $\xi$ is contained in the submodule of $Z_{G}$ generated by $Z_{G}^{k}$. Take $\varphi\in\mathopen{\hbox{${\langle}$}\kern-1.94444pt\leavevmode\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\leavevmode\hbox{${\rangle}$}}$⟩ so that $\textup{im}\varphi=\{x\in X\mid C(x)\neq\varnothing\}$ and $\varphi^{-1}x\in\partial_{\textup{ov}}^{G}C(x)$ for every $x\in\textup{im}\varphi$. Then we have $\xi\alpha_{\varphi}\in Z_{G}^{k}$ since $(\xi\alpha_{\varphi})(\cdot,x)=1_{C(\varphi x)}$ and $\partial_{\textup{ov}}^{G}C(\varphi x)\subset B_{G}(k;x)$ for $x\in\textup{dom}\varphi$. Now $\xi=\xi\alpha_{\varphi}\alpha_{\varphi^{-1}}$ holds, and thus the claim is proved.
(ii) Suppose that $\textup{H}^{1}(G,R_{G})=0$. For $k\in\mathbb{Z}_{\geq 0}$, let $[x]_{G}\setminus B_{G}(k;x)=\bigsqcup_{i=1}^{n}C_{i}(x)$ as in the proof of Corollary 4.20. Define $\xi_{i}\in Z_{G}$ by $\xi_{i}(\cdot,x)=1_{C_{i}(x)}$. Then we have $\xi_{i}\in\iota(l^{\infty}_{R}(X))+R_{G}$ for every $i$ by assumption. This implies that there exists $r>0$ such that for all $x\in X$ and $1\leq i\leq n$, either $\textup{diam}_{G}(C_{i}(x))\leq r$ or $\textup{diam}_{G}\left(\overline{C_{i}(x)}\right)\leq r$ holds. Hence $G$ is uniformly at most one-ended.
@@
The next lemma will be used to show that the above properties are preserved under some operations. Recall Definitions 1.4 and 1.5, and Notation 2.12.
@@
Let $\eta\in Z_{H}$ be defined by $\eta(\cdot,x)=1_{D(x)}$ with $D(x)\in\textup{Cut}(H)\cup\{\varnothing\}$. We define $\widehat{\eta}\in Z_{G}$ by $\widehat{\eta}(\cdot,x)=1_{\widehat{D(x)}}$, where $\widehat{\varnothing}=\varnothing$. Since $\textup{H}^{1}(G,R_{G})$ is generated by $\mathcal{C}$, there exist $\{\xi_{i}\}_{i=1}^{n}\subset Z_{G}$ satisfying $\xi_{i}(\cdot,x)=1_{C_{i}(x)}$ with $C_{i}(x)\in\mathcal{C}\cup\{\varnothing\}$, and $\{f_{i}\}_{i=1}^{n}\in l^{\infty}_{R}(X)$ such that
@@
Now we show the converse of Theorem 1.3:
@@
For $t\in T$, there exists a unique cut $C_{t}$ of $G^{\prime}$ such that $\partial_{\textup{oe}}^{G^{\prime}}C_{t}=\{t\}$ since $T$ is acyclic and $G^{\prime}=T\ast H$. We will show that $\textup{H}^{1}(G^{\prime},R_{G^{\prime}})$ is generated by the cutset $\mathcal{C}=\{C_{t}\mid t\in T\}$. Let $\xi\in Z_{G^{\prime}}$ be defined by $\xi(\cdot,y)=1_{C(y)}$ with $C(y)\in\textup{Cut}(G^{\prime})\cup\{\varnothing\}$. Since $H$ is uniformly at most one-ended, there exist $r>0$ such that if $C(y)$ intersects an $E_{H}$-class $\omega$ non-trivially (i.e., $C(y)\cap\omega\neq\varnothing,\omega$), then either $\textup{diam}_{H}(\omega\cap C(y))\leq r$ or $\textup{diam}_{H}(\omega\setminus C(y))\leq r$ holds. Then we remove the subset $\omega\cap C(y)$ from $C(y)$ if the former holds, and add the subset $\omega\setminus C(y)$ to $C(y)$ if the latter holds. This operation changes the function $\xi$ only up to $R_{G^{\prime}}$ since we only need to consider $E_{H}$-classes $\omega$ such that $H|_{\omega}\cap\partial_{\textup{ie}}^{G^{\prime}}C(y)\neq\varnothing$. Hence we may assume that for every $y\in Y$, the set $C(y)$ is $E_{H}$-invariant, which implies $\partial_{\textup{ie}}^{G^{\prime}}C(y)\subset T$. Moreover, since $C(y)$ has finitely many $G^{\prime}$-connected components (and the number of them is uniformly bounded), we may assume that every $C(y)$ is $G^{\prime}$-connected. Then we have $[y]_{G^{\prime}}\setminus C(y)=\bigsqcup_{t\in\partial_{\textup{ie}}^{G^{\prime}}C(y)}C_{t}$ since $T$ is acyclic and $G^{\prime}=T\ast H$. Hence $\xi$ is represented by $\mathcal{C}$, and $\textup{H}^{1}(G^{\prime},R_{G^{\prime}})$ is generated by $\mathcal{C}$.
@@
$C_{i}^{\prime}\subset B_{G}(r;C_{i})\textup{ and }\overline{C_{i}^{\prime}}\subset B_{G}\left(r;\overline{C_{i}}\right).$
@@
5 Proof of the theorems
5.1 Proof of Theorem 1.3
Let $R$ be a non-zero commutative ring and $(X,G)$ a Borel graph with uniformly bounded degrees. The following lemma is a key for the proof of Theorem 1.3.

### v3

@@ 4.3. Structure of $\textup{H}^{1}(G,R_{G})$
(4.5) $\sigma(a)(x)=\sum_{y\in[x]_{G}}a(x,y)\sigma(1_{\{(x,y)\}})(x).$
@@
$\sigma(a)(x)$ $=(1_{\{x\}}\sigma(a))(x)=\sigma(1_{\{x\}}a)(x)$
@@
$\Psi_{\sigma}(y,x)=\sigma(1_{\{(x,y)\}})(x)$
@@
$\Phi_{\Psi_{\sigma}}(a)(x)=\sum_{y\in[x]_{G}}a(x,y)\Psi_{\sigma}(y,x)=\sum_{y\in[x]_{G}}a(x,y)\sigma(1_{\{(x,y)\}})(x)=\sigma(a)(x),$
@@
(4.6) $\cdots\to P_{n}\to P_{n-1}\to\cdots\to P_{0}\to l^{\infty}_{R}(X)\to 0.$
@@
In other words, the set $Z_{G}^{k}$ consists of $\xi\in l^{\infty}_{R}(E_{G})$ such that the function $\xi(\cdot,x)$ is constant on each $G$-connected component of $x]_{G}\setminus B_{G}(k;x)$. Set $Z_{G}=\bigcup_{k=0}^{\infty}Z_{G}^{k}$. For $\xi\in Z_{G}$ and $a\in R_{G}$, define
@@
This function belongs to $Z_{G}$. Indeed, if $\xi\in Z_{G}^{k}$ and $\varphi\in\mathopen{\hbox{${\langle}$}\kern-1.94444pt\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\hbox{${\rangle}$}}$ with $\textup{graph}\varphi\subset G^{l}$, then we have $(\xi\alpha_{\varphi})(y,x)=\xi(y,\varphi x)1_{\textup{dom}\varphi}(x)$, and thus $\xi\alpha_{\varphi}\in Z_{G}^{k+l}$.
@@
Take $\{\varphi_{i}\}_{i=1}^{n}\subset\mathopen{\hbox{${\langle}$}\kern-1.94444pt\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\hbox{${\rangle}$}}$ so that $G=\bigsqcup_{i=1}^{n}\left(\textup{graph}\varphi_{i}\sqcup\textup{graph}\varphi_{i}^{-1}\right)$. Then by Corollary 4.7, $\textup{H}^{0}(G,W_{G})$ is identified with the right $R_{G}$-submodule
@@
$((\alpha_{\varphi_{i}}-1_{\textup{im}\varphi_{i}})\xi)(y,x)=1_{\textup{im}\varphi_{i}}(y)(\xi(\varphi_{i}^{-1}y,x)-\xi(y,x)).$
@@
$y\in[x]_{G}\mapsto 1_{\textup{im}\varphi_{i}}(y)(\xi(\varphi_{i}^{-1}y,x)-\xi(y,x))$
@@
$0$ $\rightarrow\textup{H}^{0}(G,R_{G})\rightarrow\textup{H}^{0}(G,W_{G})\xrightarrow{\pi}\textup{H}^{0}(G,W_{G}/R_{G})$
@@
It suffices to show that for every $k\in\mathbb{Z}_{\geq 0}$, the set $Z_{G}^{k}$ is contained in a finitely generated submodule of $Z_{G}$. Fix $k\in\mathbb{Z}_{\geq 0}$. For $x\in X$, let $\{C_{i}(x)\}_{i=1}^{n}$ be a family of $G$-connected components of $x]_{G}\setminus B_{G}(k;x)$ such that $x]_{G}\setminus B_{G}(k;x)=\bigsqcup_{i=1}^{n}C_{i}(x)$. By allowing $C_{i}(x)=\varnothing$, we may assume that the number $n$ does not depend on $x$. Now set $\xi_{i}(\cdot,x)=1_{C_{i}(x)}$ for $x\in X$ and $1\leq i\leq n$. Then for every $\xi\in Z_{G}^{k}$, there exists $\{f_{i}\}_{i=1}^{n}\subset l^{\infty}_{R}(X)$ such that $\xi-\sum_{i=1}^{n}\xi_{i}f_{i}\in R_{G}^{k}$. Hence $Z_{G}^{k}$ is contained in the submodule of $Z_{G}$ genetated by $\{\xi_{i}\}_{i=1}^{n}\cup\{1_{X}\}$. ∎
4.4. Cutsets and $\textup{H}^{1}(G,R_{G})$
@@
(i) Suppose that $\textup{H}^{1}(G,R_{G})$ is finitely generated as a right $R_{G}$-module. Then by Corollary 4.20, $Z_{G}$ is generated by $Z_{G}^{k}$ for some $k$. Let $x]_{G}\setminus B_{G}(k;x)=\bigsqcup_{i=1}^{n}C_{i}(x)$ as in the proof of Corollary 4.20. Then every element of $Z_{G}^{k}$ is represented by the cutset
@@
which has uniformly bounded boundaries. Note that if $\xi\in Z_{G}$ satisfies $\xi(\cdot,x)=1_{C(x)}$, then for $\varphi\in\mathopen{\hbox{${\langle}$}\kern-1.94444pt\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\hbox{${\rangle}$}}$, we have $(\xi\alpha_{\varphi})(\cdot,x)=1_{C(\varphi x)}1_{\textup{dom}\varphi}(x)$. Hence every element of $Z_{G}$ is represented by the cutset $\mathcal{C}$.
@@
Let $\xi\in Z_{G}$ be such that $\xi(\cdot,x)=1_{C(x)}$ with $C(x)\in\mathcal{C}\cup\{\varnothing\}$. It suffices to show that $\xi$ is contained in the submodule of $Z_{G}$ generated by $Z_{G}^{k}$. Take $\varphi\in\mathopen{\hbox{${\langle}$}\kern-1.94444pt\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\hbox{${\rangle}$}}$ so that $\textup{im}\varphi=\{x\in X\mid C(x)\neq\varnothing\}$ and $\varphi^{-1}x\in\partial_{\textup{ov}}^{G}C(x)$ for every $x\in\textup{im}\varphi$. Then we have $\xi\alpha_{\varphi}\in Z_{G}^{k}$ since $(\xi\alpha_{\varphi})(\cdot,x)=1_{C(\varphi x)}$ and $\partial_{\textup{ov}}^{G}C(\varphi x)\subset B_{G}(k;x)$ for $x\in\textup{dom}\varphi$. Now $\xi=\xi\alpha_{\varphi}\alpha_{\varphi^{-1}}$ holds, and thus the claim is proved.
(ii) Suppose that $\textup{H}^{1}(G,R_{G})=0$. For $k\in\mathbb{Z}_{\geq 0}$, let $x]_{G}\setminus B_{G}(k;x)=\bigsqcup_{i=1}^{n}C_{i}(x)$ as in the proof of Corollary 4.20. Define $\xi_{i}\in Z_{G}$ by $\xi_{i}(\cdot,x)=1_{C_{i}(x)}$. Then we have $\xi_{i}\in\iota(l^{\infty}_{R}(X))+R_{G}$ for every $i$ by assumption. This implies that there exists $r>0$ such that for all $x\in X$ and $1\leq i\leq n$, either $\textup{diam}_{G}(C_{i}(x))\leq r$ or $\textup{diam}_{G}\left(\overline{C_{i}(x)}\right)\leq r$ holds. Hence $G$ is uniformly at most one-ended.
@@
The next lemma will be needed to show that the above properties are preserved after decomposing a graph. Recall Definitions 1.4 and 1.5, and Notation 2.12.
@@
Let $\eta\in Z_{H}$ be defined by $\eta(\cdot,x)=1_{D(x)}$ with $D(x)\in\textup{Cut}(H)\cup\{\varnothing\}$. We define $\widehat{\eta}\in Z_{G}$ by $\widehat{\eta}(\cdot,x)=1_{\widehat{D(x)}}$, where $\widehat{\varnothing}$$\varnothing$. Since $\textup{H}^{1}(G,R_{G})$ is generated by $\mathcal{C}$, there exist $\{\xi_{i}\}_{i=1}^{n}\subset Z_{G}$ satisfying $\xi_{i}(\cdot,x)=1_{C_{i}(x)}$ with $C_{i}(x)\in\mathcal{C}\cup\{\varnothing\}$, and $\{f_{i}\}_{i=1}^{n}\in l^{\infty}_{R}(X)$ such that
@@
Now we show the converse of Theorem A:
@@
For $t\in T$, there exists a unique cut $C_{t}$ of $G^{\prime}$ such that $\partial_{\textup{oe}}^{G^{\prime}}C_{t}=\{t\}$ since $T$ is acyclic and $G^{\prime}=T\ast H$. We will show that $\textup{H}^{1}(G^{\prime},R_{G^{\prime}})$ is generated by the cutset $\mathcal{C}=\{C_{t}\mid t\in T\}$. Let $\xi\in Z_{G^{\prime}}$ be defined by $\xi(\cdot,y)=1_{C(y)}$ with $C(y)\in\textup{Cut}(G^{\prime})\cup\{\varnothing\}$. Since $H$ is uniformly at most one-ended, there exist $r>0$ such that if $C(y)$ intersects an $E_{H}$-class $\omega$ non-trivially (i.e., $C(y)\cap\omega\neq\varnothing,\omega$), then either $\textup{diam}_{H}(\omega\cap C(y))\leq r$ or $\textup{diam}_{H}(\omega\setminus C(y))\leq r$ holds. Then we remove the subset $\omega\cap C(y)$ from $C(y)$ if the former holds, and add the subset $\omega\setminus C(y)$ to $C(y)$ if the latter holds. This operation changes the function $\xi$ only up to $R_{G^{\prime}}$ since we only need to consider $E_{H}$-classes $\omega$ such that $H|_{\omega}\cap\partial_{\textup{ie}}^{G^{\prime}}C(y)\neq\varnothing$. Hence we may assume that for every $y\in Y$, the set $C(y)$ is $E_{H}$-invariant, which implies $\partial_{\textup{ie}}^{G^{\prime}}C(y)\subset T$. Moreover, since $C(y)$ has finitely many $G^{\prime}$-connected components (and the number of them is uniformly bounded), we may assume that every $C(y)$ is $G^{\prime}$-connected. Then we have $y]_{G^{\prime}}\setminus C(y)=\bigsqcup_{t\in\partial_{\textup{ie}}^{G^{\prime}}C(y)}C_{t}$ since $T$ is acyclic and $G^{\prime}=T\ast H$. Hence $\xi$ is represented by $\mathcal{C}$, and $\textup{H}^{1}(G^{\prime},R_{G^{\prime}})$ is generated by $\mathcal{C}$.
@@
[…]
@@
5. Proof of the theorems
5.1. Proof of Theorem A
Let $R$ be a non-zero commutative ring and $(X,G)$ a Borel graph with uniformly bounded degrees. The following lemma is a key for the proof of Theorem A.